Find the area or region that corresponds to each of the following z values. Sketch



and label each curve.



1. to the right of z = 2.18



2. below z= -1.40



3. above z= 0.57



4. between z = -2.34 and z = 1.57



5. between z = 0.32 and z = 2.42

1
Expert's answer
2022-03-08T05:25:07-0500

1) "P(Z>2.18)=1-P(Z<2.18)=0.0146."

2) "P(Z<-1.40)=0.0808."

3) "P(Z>0.57)=1-P(Z<0.57)=0.2843."

4) "P(-2.34<Z<1.57)=P(Z<1.57)-P(Z<-2.34)=0.9322."

5) "P(0.32<Z<2.42)=P(Z<2.42)-P(Z<0.32)=0.3667."